OpenStudy (anonymous):

class has 10 girls and 16 boys. How many boy-girl dates can be arranged?

OpenStudy (kinggeorge):

Let's look at the choices the girls can have. The first girl has 16 choices, the second girl has 15, the third 14, and so on until we get to the last girl with 1 choice. We need to multiply all these together to get our answer. Thus, we get \[16\cdot15\cdot14\cdot ... \cdot 8\cdot7\]possibilities.

OpenStudy (kinggeorge):

If you want an actual number, it turns out to be 29059430400

OpenStudy (anonymous):

thanks!

OpenStudy (kinggeorge):

you're welcome

OpenStudy (anonymous):

Thats alot of dates. I thought you would just multiply 10 x 16 and get 160 dates.

OpenStudy (kinggeorge):

Nope. You have to use permutations here.

OpenStudy (anonymous):

so, do you have a formula for permutations?

OpenStudy (kinggeorge):

\[n\text{P}k={n! \over (n-k)!}\]In this case, we have \(n=16\) and \(k=10\)

OpenStudy (anonymous):

thanks again!